For which function is given below, there are is a maximum and a minimum values on the given interval?

a) f(x)=x2 on (0,1)

b) f(x)=1/x if x=0 on [0,1] , 0 if x=0

c) f(x)=2x+1 on R

d) f(x)=x2+1/x on [1,2]

e) None of the above.


1
Expert's answer
2021-02-04T08:02:45-0500

a) No. "f(x)=x^2" increases on open interval "(0, 1)" and is undefined at "x=0" and "x=1."


b) No. "f(x)=1\/x" decreases on open interval "(0, 1)" from "\infin" to "1."

"f(x)" has minimum value on "[0, 1]," but has no maximum value on "[0, 1]."


c) No. "f(x)=2x+1" increases from "-\infin" to "\infin" on "\R."


d) Yes. The function "f(x)=x^2+1\/x" is defined on "[1, 2]." "f'(x)=2x-1\/x^2"

"f'(x)>0" on "[1, 2]" => "f(x)" increases on "(1, 2)"

"f(1)=2, f(2)=9\/2"

The function "f(x)=x^2+1\/x" has the maximum with value of "9\/2" on "[1, 2]" at "x=2."

The function "f(x)=x^2+1\/x" has the minimum with value of "2" on "[1, 2]" at "x=1."


d) "f(x)=x^2+1\/x" on "[1, 2]."